# Relative equilibria for Scaling Symmetries and Central Configurations

### v1

@@ 4.1 Scaled Cotangent Lifted Action
Given an action of $G=\mathbb{R}^{+}$ on a manifold $Q$, we can define a corresponding action on the cotangent bundle $T^{*}Q$. The standard method for this is using the cotangent lift [17]. Here, however, we define a “scaled” cotangent lift, which is more general.
@@
Let $\Psi$ be an action of the group $G=\mathbb{R}^{+}$ on a manifold $Q$. For any $g\in G$, the scaled cotangent lift $\hat{\Psi}_{g}^{c}$ with parameter $c\in\mathbb{R}$ of the diffeomorphism $\Psi_{g}$ to $T^{*}Q$ is defined as:
$\left\langle\hat{\Psi}_{g}^{c}(\alpha),v\right\rangle=\left\langle\alpha,g^{c}(T\Psi_{g^{-1}}\cdot v)\right\rangle,$
where $\alpha\in T^{*}_{q}Q$ is a cotangent vector, $v\in T_{\Psi_{g}(q)}Q$ is a tangent vector, and $\Psi_{g^{-1}}=(\Psi_{g})^{-1}$. Then the scaled cotangent lifted action with parameter $c$ is the action $\hat{\Psi}^{c}:G\times T^{*}Q\to T^{*}Q$ given by
$\hat{\Psi}^{c}(g,(q,\alpha))=\hat{\Psi}_{g}^{c}(q,\alpha).$ (4)
If $\Psi$ is the action of $G=\mathbb{R}^{+}$ on $Q$ in local coordinates $(\mathbf{q},\mathbf{p})$ on $T^{\ast}Q$ , we can express the scaled cotangent lifted action with parameter $c$ on $T^{\ast}Q$ as
$\hat{\Psi}^{c}(g,(\mathbf{q},\mathbf{p}))=\left(g,\Psi_{g}(\mathbf{b}),g^{c}(D\Psi_{g}(\mathbf{q}))^{-T}\cdot\mathbf{p}\right).$ (5)
@@
Every scaled cotangent lifted action of $G=\mathbb{R}^{+}$ is conformally symplectic with respect to the canonical symplectic form $\omega_{0}$.
@@
Let $G=\mathbb{R}^{+}$ act on $Q$ via the action $\Psi^{c}:G\times Q\to Q$. Let $\hat{\Psi}^{c}:G\times T^{\ast}Q\to T^{\ast}Q$ denote the scaled cotangent lifted action on $T^{\ast}Q$ with parameter $c$. Let $\xi\in\mathfrak{g}=\mathbb{R}$ be an element of the Lie algebra of $G$, and let $\xi_{Q}$ be the infinitesimal generator of the action $\Psi$ on $Q$ associated with $\xi$. Then, in local coordinates, the infinitesimal generator of the action $\hat{\Psi}^{c}$ on $T^{\ast}Q$ is given by
@@
Let $\Psi$ be an action of $G$ on $Q$, and let $g=\exp t\xi$. Then by definition
$\xi_{Q}(\mathbf{q})=\left.\frac{d}{dt}\right|_{t=0}\Psi_{\exp t\xi}(\mathbf{q}).$
Therefore, using equation (5)
@@
$\left\langle(D\Psi_{\exp t\xi}(\mathbf{q}))^{-T}\cdot\mathbf{p},D\Psi_{\exp t\xi}(\mathbf{q})\cdot\mathbf{v}\right\rangle=\left\langle\mathbf{v},\mathbf{p}\right\rangle$
for all $t$, differentiating with respect to $t$ gives,
@@
$=-\left\langle(D\xi_{Q}(\mathbf{q}))^{T}\cdot\mathbf{p},\mathbf{v}\right\rangle,$
which implies that $\left.\frac{d}{dt}\right|_{t=0}(D\Psi_{\exp t\xi}(\mathbf{q}))^{-T}\cdot\mathbf{p}=-(D\xi_{Q}(\mathbf{q}))^{T}\cdot\mathbf{p}$ since $\mathbf{v}$ is arbitrary. ∎
@@ 5 Conformal Momentum Map
Let $(M,\omega,\theta)$ an exact symplectic manifold and $\Phi:G\times M\to M$ the action of the multiplicative group $G=\mathbb{R}^{+}$ on $M$. Suppose $\Phi$ is a conformally symplectic action with parameter $c$. We say that a map $J:M\to\mathfrak{g}^{\ast}=\mathbb{R}$ is a conformal momentum map with parameter $c$ for $\Phi$ provided that for every $\xi\in\mathfrak{g}=\mathbb{R}$
@@
Let $(M,\omega,\theta)$ an exact symplectic manifold. Let $\Phi$ be a conformally symplectic action such that $\Phi_{g}^{\ast}\theta=g^{c}\theta$ for all $g\in G=\mathbb{R}^{+}$. Then $J:M\to\mathfrak{g}^{\ast}=\mathbb{R}^{+}$ defined by
@@ 5.1 Conformal Momentum Maps for Actions on Cotangent Bundles
Suppose $Q$ is a manifold and $M=T^{\ast}Q$ is the cotangent bundle with the canonical forms. Let $\Psi$ be an action of $G=\mathbb{R}^{+}$, and let $\hat{\Psi}^{c}$ be the scaled cotangent lifted action with parameter $c$ of $\Psi$ to $T^{\ast}Q$. The following theorem provides a specialized formula for the conformal momentum map in this context, which will be useful in examples and applications.
@@
Let $\Psi$ be an action of $G=\mathbb{R}^{+}$ on $Q$, and let $\xi\in\mathfrak{g}=\mathbb{R}$. Suppose $\hat{\Psi}^{c}$ is the scaled cotangent lifted action with parameter $c$ of $\Psi$ Then,the conformal momentum map $J:T^{\ast}Q\to\mathfrak{g}^{\ast}=\mathbb{R}$ is obtained by the formula:
$J_{\xi}(p_{q})=\left\langle J(p_{q}),\xi\right\rangle=\left\langle p_{q},\xi_{Q}(q)\right\rangle,$ (8)
where $p_{q}\in T^{\ast}Q$, and $\xi_{Q}$ is the infinitesimal generator of $\Psi$ on $Q$. In local coordinates $(\mathbf{q},\mathbf{p})$, the formula takes the form
$J_{\xi}(\mathbf{q},\mathbf{p})=\xi J(\mathbf{q},\mathbf{p})=\left\langle\mathbf{p},\xi_{Q}(\mathbf{q})\right\rangle,$ (9)
where for every $\xi\in\mathfrak{g}$ the conformal momentum function $J_{\xi}$ satisfies the equation $J_{\xi}(\mathbf{q},\mathbf{p})=\xi J(\mathbf{q},\mathbf{p})$.
@@
The following example will be useful when discussing the $N$-body problem.
@@
Let $Q=\mathbb{R}^{3n}$ and its cotangent bundle $T^{\ast}Q$ with local coordinates $(\mathbf{q},\mathbf{p})\in T^{\ast}Q$. Let $\Psi$ the group action of the group $G=\mathbb{R}^{+}$ on $Q=\mathbb{R}^{3n}$ defined by
$\Psi(g,\mathbf{q})=(g,g\mathbf{q})$
for all $g\in\mathbb{R}^{+}$. Consider the scaled cotangent lifted action $\hat{\Psi}^{c}$ defined by
$\hat{\Psi}^{c}(g,(\mathbf{q},\mathbf{p}))$ $=\left(g,\Psi_{g}(\mathbf{q}),g^{c}(D\Psi_{g}(\mathbf{q}))^{-T}\mathbf{p}\right)$
$=\left(g,g\mathbf{q},g^{c-1}\mathbf{p}\right)$
The choice $c=1/2$ gives the Kepler scalings, which will use in other examples. Clearly,
$\xi_{Q}(\mathbf{q})=\left.\frac{d}{dt}\right|_{t=0}\Psi_{\exp(t\xi)}=\left.\frac{d}{dt}\right|_{t=0}(e^{t\xi})\mathbf{q}=\xi\mathbf{q}.$
Consequently, considering $\mathbf{p}$ as a column vector, the conformal momentum function associated to $\xi$ is $J_{\xi}(\mathbf{q},\mathbf{p})=\left\langle\mathbf{p},\xi_{Q}(\mathbf{q})\right\rangle=\xi\mathbf{p}^{T}\mathbf{q}$, and the momentum map is $J(\mathbf{q},\mathbf{p})=\mathbf{p}^{T}\mathbf{q}$.
As a simple exercise we can easily compute $\hat{\xi}_{T^{\ast}Q}^{c}(\mathbf{q})$ using equation (6). Explicitely, we have
$\hat{\xi}^{c}_{T^{\ast}Q}(\mathbf{q})$ $=\left(\xi_{Q}(\mathbf{q}),\left((c\xi)\operatorname{Id}-\left(D\xi_{Q}(\mathbf{q})\right)^{T}\right)\mathbf{p}\right)$
$=(\xi\mathbf{q},(c-1)\xi\mathbf{p})$
$=\left(\xi\mathbf{q},-\frac{1}{2}\xi\mathbf{p}\right).$
As another simple computation we can verify that $J$ is indeed a conformal momentum map with parameter $c=1/2$, by verifying that $\hat{\xi}_{T^{\ast}Q}^{c}$ is a conformally Hamiltonian vector field with conformal Hamiltonian $J_{\xi}$. Explicitely, we have that
$\frac{\partial J_{\xi}}{\partial\mathbf{p}}(\mathbf{q},\mathbf{p})=\xi\mathbf{q}=\xi_{Q}(\mathbf{q})$
and if $c=1/2$, then
$-\frac{\partial J_{\xi}}{\partial\mathbf{q}}+(c\xi)\mathbf{p}=-\frac{\partial J_{\xi}}{\partial\mathbf{q}}+\xi\mathbf{p}=-\xi\mathbf{p}+\frac{1}{2}\xi\mathbf{p}=-\frac{1}{2}\xi\mathbf{p}$
Hence,
$\left(\frac{\partial J_{\xi}}{\partial\mathbf{p}},-\frac{\partial J_{\xi}}{\partial\mathbf{q}}+\frac{1}{2}\xi\mathbf{p}\right)=\hat{\xi}_{T^{\ast}Q}^{c},$
which confirms that $J$ is a conformal momentum map, since $J_{\xi}$ is a conformal Hamiltonian with parameter $c=1/2$ for the conformally Hamiltonian vector field $\hat{\xi}_{T^{\ast}Q}^{c}$.
@@ 5.2 Properties of Conformal Momentum Maps
Let $(M,\omega,\theta)$ be an exact symplectic manifold. Consider $\Phi$ to be a conformally symplectic action with parameter $c$ of $G=\mathbb{R}^{+}$ on $M$ with a conformal momentum map $J$. If $H$ is conformally invariant with parameter $b$, then along the flow of $X_{H}$, the momentum map $J$ satisfies the equation:
@@
We now present another property of the conformal momentum map: under certain hypotheses, the momentum map is a scaling function. This implies that the conformal momentum map can be employed in the reduction studied in [5]. We begin with the following lemma, which is essentially an infinitesimal version of Theorem 5.3:
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold. Suppose that $F:M\to\mathbb{R}$ is a smooth function and $X_{F}^{c}$ is a conformal vector field with parameter $c$, associated with $F$. Then, the following statements hold:
@@
    Conversely, if $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$, then $F=i_{X_{F}^{c}}\theta+\text{constant}$.
@@
Suppose $F=i_{X_{F}^{c}}\theta$. Since $dF=c\theta+i_{X_{F}^{c}}\omega$ we have that $d(i_{X_{F}^{c}}\theta)=c\theta-i_{X_{F}^{c}}d\theta$, and by Cartan’s magic formula
$c\theta=d(i_{X_{F}^{c}}\theta)+i_{X_{F}^{c}}d\theta=\mathcal{L}_{X_{F}^{c}}\theta.$
Conversely if $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$, by Cartan’s magic formula $c\theta=d(i_{X_{F}^{c}}\theta)+i_{X_{F}^{c}}d\theta=\mathcal{L}_{X_{F}^{c}}\theta$ and since $dF=c\theta+i_{X_{F}^{c}}\omega$ we find
@@
that is, $F=i_{X_{F}^{c}}\theta+constant$. ∎
@@
Let $(M,\omega,\theta)$ an exact symplectic manifold, $F:M\to\mathbb{R}$ a smooth function and $X_{F}^{c}$ be a conformally Hamiltonian vector field with parameter $c$, associated to $F$. If $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$ then the function $F$ is a scaling function with parameter $c$ for $X_{F}^{c}$.
@@
Suppose $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$. Then, by Lemma 5.7, we have $F=i_{X_{F}^{c}}\theta+\text{constant}$. Consequently, we obtain:
$\mathcal{L}_{X_{F}^{c}}F$ $=i_{X_{F}^{c}}dF$
$=i_{X_{F}^{c}}(d(i_{X_{F}^{c}}\theta))$
$=i_{X_{F}^{c}}(\mathcal{L}_{X_{F}^{c}}\theta-i_{X_{F}^{c}}d\theta)$
@@
$=cF.$
Therefore, $\mathcal{L}_{X_{F}^{c}}F=cF$, as required. ∎
Corollary 5.9.
Let $(M,\omega,\theta)$ be an exact symplectic manifold, and let $J:M\to\mathbb{R}$ be the conformal momentum map associated with a conformally symplectic action of the Lie group $G=\mathbb{R}^{+}$. Suppose $X_{J}^{c}$ is the conformally Hamiltonian vector field with parameter $c$ associated with $J$. If $\mathcal{L}_{X_{J}^{c}}\theta=c\theta$, then the conformal momentum map $J$ is a scaling function with parameter $c$ for $X_{J}^{c}$.
Proof.
Since the momentum map $J$ satisfies the equation $i_{X_{J}^{c}}\omega+c\theta=dJ$, where $X_{J}^{c}$ is the infinitesimal generator corresponding to the Lie algebra element $1$, it follows immediately from Proposition 5.8 that $J$ is a scaling function with parameter $c$ for $X_{J}^{c}$. ∎
Suppose $G$ is a group acting on $Q$, with an induced action on $T^{*}Q$ by cotangent lift. It is well known that the cotangent lifted action admits an equivariant momentum map, which is given by the formula
$\left\langle J(\alpha_{q}),\xi\right\rangle=\left\langle\alpha_{q},\xi_{Q}(q)\right\rangle,$
where $\alpha_{q}=(q,\alpha)\in T^{*}Q$. In the Abelian case, the concept of equivariance simplifies to invariance due to the triviality of the adjoint action. The equivariance of the momentum map is crucial because it ensures the compatibility of the momentum map with the group action, which is necessary for performing symplectic reduction.
Unfortunately, the conformal momentum map associated with the scaled cotangent lifted action, as described in equation (8), is not equivariant. Nevertheless, we have previously shown that the scaled cotangent lifted action is infinitesimally conformally invariant. As we will demonstrate below, it is, in fact, fully conformally invariant. We begin with a definition that generalizes the concept of equivariance.
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold with a conformally symplectic action $\Phi$ (with parameter $c$) of the group $G=\mathbb{R}^{+}$ on $M$, and let $J$ be a conformal momentum map for $\Phi$. Then $J$ is said to be conformally equivariant if for all $g\in G$,
@@
where $\Psi$ is the action of $G$ on $Q$.
Since $G=\mathbb{R}^{+}$, the adjoint action is trivial. Hence,

### v2

@@ 4.1 Scaled Cotangent Lifted Action
Given an action of the multiplicative group $\mathbb{R}^{+}$ on a manifold $Q$, we can define a corresponding action on the cotangent bundle $T^{*}Q$. The standard method for this is using the cotangent lift [16]. Here, however, we define a “scaled” cotangent lift, which is more general.
@@
Let $\Psi$ be an action of the group $\mathbb{R}^{+}$ on a manifold $Q$. For any $g\in\mathbb{R}^{+}$ and $z=(p,q)\in T^{\ast}Q$ , the scaled cotangent lift $\hat{\Psi}_{g}^{c}$ with parameter $c\in\mathbb{R}$ of the diffeomorphism $\Psi_{g}$ to $T^{*}Q$ is defined as:
$\hat{\Psi}_{g}^{c}(z)=(\Psi_{g}(q),g^{c}\cdot(d\Psi_{g^{-1}})^{\ast}(p))$
where $\Psi_{g^{-1}}=(\Psi_{g})^{-1}$. Then the scaled cotangent lifted action with parameter $c$ is the action $\hat{\Psi}^{c}:\mathbb{R}^{+}\times T^{*}Q\to T^{*}Q$ given by
$\hat{\Psi}^{c}(g,z)=\hat{\Psi}_{g}^{c}(z).$ (4)
If $\Psi$ is the action of $\mathbb{R}^{+}$ on $Q$ in local coordinates $(\mathbf{q},\mathbf{p})$ on $T^{\ast}Q$ , we can express the scaled cotangent lifted action with parameter $c$ on $T^{\ast}Q$ as
$\hat{\Psi}^{c}(g,(\mathbf{q},\mathbf{p}))=\left(\Psi_{g}(\mathbf{q}),g^{c}(D\Psi_{g}(\mathbf{q}))^{-T}\cdot\mathbf{p}\right).$ (5)
@@
Every scaled cotangent lifted action of the multiplicative group $\mathbb{R}^{+}$ is conformally symplectic with respect to the canonical symplectic form $\omega_{0}$.
@@
Let $\mathbb{R}^{+}$ act on $Q$ via the action $\Psi^{c}:\mathbb{R}^{+}\times Q\to Q$. Let $\hat{\Psi}^{c}:\mathbb{R}^{+}\times T^{\ast}Q\to T^{\ast}Q$ denote the scaled cotangent lifted action on $T^{\ast}Q$ with parameter $c$. Let $\xi\in\mathbb{R}$ be an element of the Lie algebra of $\mathbb{R}^{+}$, and let $\xi_{Q}$ be the infinitesimal generator of the action $\Psi$ on $Q$ associated with $\xi$. Then, in local coordinates, the infinitesimal generator of the action $\hat{\Psi}^{c}$ on $T^{\ast}Q$ is given by
@@
Let $\Psi$ be an action of $\mathbb{R}^{+}$ on $Q$, and define $g=\exp(t\xi)$. By definition, the infinitesimal generator $\xi_{Q}$ is given by:
$\xi_{Q}(\mathbf{q})=\left.\frac{d}{dt}\right|_{t=0}\Psi_{\exp(t\xi)}(\mathbf{q}).$
Now, using equation (5), we compute the infinitesimal generator $\hat{\xi}_{T^{*}Q}^{c}$ of the cotangent lift:
@@
$\left\langle(D\Psi_{\exp t\xi}(\mathbf{q}))^{-T}\cdot\mathbf{p},D\Psi_{\exp t\xi}(\mathbf{q})\cdot\mathbf{v}\right\rangle=\left\langle\mathbf{p},\mathbf{v}\right\rangle$
for all $t$, differentiating both sides with respect to $t$ gives,
@@
$=-\left\langle(D\xi_{Q}(\mathbf{q}))^{T}\cdot\mathbf{p},\mathbf{v}\right\rangle.$
Since $\mathbf{v}$ is arbitrary, we conclude:
$\left.\frac{d}{dt}\right|_{t=0}(D\Psi_{\exp(t\xi)}(\mathbf{q}))^{-T}\cdot\mathbf{p}=-(D\xi_{Q}(\mathbf{q}))^{T}\cdot\mathbf{p}.$
@@ 5 Conformal Momentum Map
Let $(M,\omega,\theta)$ an exact symplectic manifold and $\Phi:\mathbb{R}^{+}\times M\to M$ the action of the multiplicative group $\mathbb{R}^{+}$ on $M$. Suppose $\Phi$ is a conformally symplectic action with parameter $c$. We say that a map $J:M\to\mathbb{R}$ is a conformal momentum map with parameter $c$ for $\Phi$ provided that for every $\xi\in\mathbb{R}$
@@
Let $(M,\omega,\theta)$ an exact symplectic manifold. Let $\Phi$ be a conformally symplectic action such that $\Phi_{g}^{\ast}\theta=g^{c}\theta$ for all $g\in\mathbb{R}^{+}$. Then $J:M\to\mathbb{R}^{+}$ defined by
@@ 5.1 Conformal Momentum Maps for Actions on Cotangent Bundles
Suppose $Q$ is a manifold and $M=T^{\ast}Q$ is the cotangent bundle with the canonical forms. Let $\Psi$ be an action of the multiplicative group $\mathbb{R}^{+}$ on $Q$, and let $\hat{\Psi}^{c}$ be the scaled cotangent lifted action with parameter $c$ of $\Psi$ to $T^{\ast}Q$. The following theorem provides a specialized formula for the conformal momentum map in this context, which will be useful in examples and applications.
@@
Let $\Psi$ be an action of the multiplicative group $\mathbb{R}^{+}$ on a manifold $Q$, and let $\xi\in\mathbb{R}$ be an element of the Lie algebra of $\mathbb{R}^{+}$. Suppose $\hat{\Psi}^{c}$ is the scaled cotangent lift of $\Psi$ with parameter $c$. Then, the conformal momentum map $J:T^{*}Q\to\mathbb{R}$ is given by:
$J_{\xi}(z)=\left\langle J(z),\xi\right\rangle=\left\langle p,\xi_{Q}(q)\right\rangle,$ (8)
where $z=(q,p)\in T^{*}Q$, with $q=\pi(z)$ as the base point under the canonical projection $\pi:T^{*}Q\to Q$, and $p\in T^{*}_{q}Q$. Here, $\xi_{Q}$ denotes the infinitesimal generator of the action $\Psi$ on $Q$, and the pairing on the right-hand side is between the covector $p\in T_{q}^{*}Q$ and the vector $\xi_{Q}(q)\in T_{q}Q$.
In local coordinates $(\mathbf{q},\mathbf{p})$, the formula becomes:
$J_{\xi}(\mathbf{q},\mathbf{p})=\xi J(\mathbf{q},\mathbf{p})=\left\langle\mathbf{p},\xi_{Q}(\mathbf{q})\right\rangle.$ (9)
@@
The following example will be useful for our discussion of the Newtonian $N$-body problem.
@@
Consider the configuration space $Q=\mathbb{R}^{3n}$ and its cotangent bundle $T^{*}Q$ with local coordinates $(\mathbf{q},\mathbf{p})\in T^{*}Q$. Let $\Psi:\mathbb{R}^{+}\times\mathbb{R}^{3n}\to\mathbb{R}^{3n}$ denote the group action of the multiplicative group $\mathbb{R}^{+}$ on $Q$, defined by:
$\Psi(g,\mathbf{q})=g\mathbf{q}\quad\text{for all }g\in\mathbb{R}^{+}.$
The corresponding scaled cotangent lifted action $\hat{\Psi}^{c}$ is given by:
$\hat{\Psi}^{c}(g,(\mathbf{q},\mathbf{p}))=\left(g\mathbf{q},g^{c-1}\mathbf{p}\right).$
The choice $c=1/2$ gives the Kepler scalings, which we will is useful when studying the Newtonian $N$-body problem. Clearly, the infinitesimal generator $\xi_{Q}$ of the action on $Q$ is:
$\xi_{Q}(\mathbf{q})=\left.\frac{d}{dt}\right|_{t=0}\Psi_{\exp(t\xi)}=\xi\mathbf{q}.$
Thus, considering $\mathbf{p}$ as a column vector, the conformal momentum function associated with $\xi$ is:
$J_{\xi}(\mathbf{q},\mathbf{p})=\left\langle\mathbf{p},\xi_{Q}(\mathbf{q})\right\rangle=\xi\mathbf{p}^{T}\mathbf{q},$
and the momentum map is:
$J(\mathbf{q},\mathbf{p})=\mathbf{p}^{T}\mathbf{q}.$
As an exercise, we can compute $\hat{\xi}_{T^{*}Q}^{c}(\mathbf{q})$ using equation (6). Explicitly, we have:
$\hat{\xi}_{T^{*}Q}^{c}(\mathbf{q})=\left(\xi_{Q}(\mathbf{q}),\left(c\xi\operatorname{Id}-(D\xi_{Q}(\mathbf{q}))^{T}\right)\mathbf{p}\right)=\left(\xi\mathbf{q},(c-1)\xi\mathbf{p}\right),$
which for $c=1/2$ simplifies to:
$\hat{\xi}_{T^{*}Q}^{c}(\mathbf{q})=\left(\xi\mathbf{q},-\frac{1}{2}\xi\mathbf{p}\right).$
Next, we verify that $J$ is indeed a conformal momentum map with parameter $c=1/2$ by showing that $\hat{\xi}_{T^{*}Q}^{c}$ is a conformally Hamiltonian vector field with conformal Hamiltonian $J_{\xi}$. Explicitly, we compute:
$\frac{\partial J_{\xi}}{\partial\mathbf{p}}(\mathbf{q},\mathbf{p})=\xi\mathbf{q}=\xi_{Q}(\mathbf{q}),$
and for $c=1/2$, we have:
$-\frac{\partial J_{\xi}}{\partial\mathbf{q}}+(c\xi)\mathbf{p}=-\xi\mathbf{p}+\frac{1}{2}\xi\mathbf{p}=-\frac{1}{2}\xi\mathbf{p}.$
Thus, we confirm:
$\left(\frac{\partial J_{\xi}}{\partial\mathbf{p}},-\frac{\partial J_{\xi}}{\partial\mathbf{q}}+\frac{1}{2}\xi\mathbf{p}\right)=\hat{\xi}_{T^{*}Q}^{c},$
which shows that $J$ is a conformal momentum map, since $J_{\xi}$ is a conformal Hamiltonian with parameter $c=1/2$ for the conformally Hamiltonian vector field $\hat{\xi}_{T^{*}Q}^{c}$.
@@ 5.2 Properties of Conformal Momentum Maps
In this subsection, we explore key properties of conformal momentum maps in exact symplectic manifolds with conformally symplectic actions. We begin by generalizing Noether’s theorem to the setting of conformally invariant Hamiltonians, leading to new conservation laws for conformal momentum maps. We then explore the connection between conformal momentum maps and scaling functions, establishing conditions under which the momentum map serves as a scaling function. Finally, we extend the notion of equivariance to conformal equivariance, which in this context reduces to conformal invariance. We demonstrate how the conformal momentum map exhibits this conformal invariance in scaled cotangent lifted actions.
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold. Consider $\Phi$ to be a conformally symplectic action with parameter $c$ of $\mathbb{R}^{+}$ on $M$ with a conformal momentum map $J$. If $H$ is conformally invariant with parameter $b$, then along the flow of $X_{H}$, the momentum map $J$ satisfies the equation:
@@
We now establish an important property of conformally Hamiltonian vector fields. This result lays the foundation for understanding how conformal momentum maps relate to scaling functions:
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold. Suppose $F:M\to\mathbb{R}$ is a smooth function, and let $X_{F}^{c}$ be a conformally Hamiltonian vector field with parameter $c$, with conformal Hamiltonian $F$. Then the following statements hold:
@@
    If $M$ is connected and $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$, then $F=i_{X_{F}^{c}}\theta+\text{constant}$.
@@
Suppose $F=i_{X_{F}^{c}}\theta$. By definition of conformally Hamiltonian vector field we have:
$dF=d(i_{X_{F}^{c}}\theta)=c\theta+i_{X_{F}^{c}}\omega.$
Since $\omega=-d\theta$, using Cartan’s magic formula $\mathcal{L}_{X_{F}^{c}}\theta=d(i_{X_{F}^{c}}\theta)+i_{X_{F}^{c}}d\theta$, we obtain:
$\mathcal{L}_{X_{F}^{c}}\theta=d(i_{X_{F}^{c}}\theta)-i_{X_{F}^{c}}\omega=c\theta.$
This proves the first direction.
Conversely, suppose $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$. Using Cartan’s formula again, we have:
$c\theta=d(i_{X_{F}^{c}}\theta)-i_{X_{F}^{c}}\omega.$
Rearranging, we get:
@@
which implies $F=i_{X_{F}^{c}}\theta+\text{constant}$, since $M$ is connected, and any two functions with the same differential must differ by a constant on a connected domain. ∎
Next, we demonstrate that the conformal momentum map defined in Theorem 5.3, is, in fact, a scaling function under the conformally symplectic action. This result ties directly to the reduction framework studied in [3]:
Corollary 5.8.
Let $(M,\omega,\theta)$ be an exact symplectic manifold, and let $\Phi$ be a conformally symplectic action with parameter $c$. Consider the momentum map $J:M\to\mathbb{R}$, defined by $J_{\xi}=\xi J=i_{\xi_{M}}\theta$ as described in Theorem 5.3, where $\xi_{M}$ is the infinitesimal generator of the action corresponding to the Lie algebra element $\xi\in\mathbb{R}$. With this choice of momentum map, $J$ is a scaling function with parameter $c$.
Proof.
The conformal momentum map $J$ satisfies the equation $i_{X_{J}^{c}}\omega+c\theta=dJ$, where $X_{J}^{c}$ is the infinitesimal generator corresponding to the Lie algebra element $1$. Since $J=i_{X_{J}^{c}}\theta$, by Lemma 5.7, we know that $\mathcal{L}_{X_{J}^{c}}\theta=c\theta$.
To confirm $J$ as a scaling function, we compute:
$\mathcal{L}_{X_{J}^{c}}J$ $=i_{X_{J}^{c}}d{J}$
$=i_{X_{J}^{c}}(d(i_{X_{J}^{c}}\theta))\quad(\text{since }J=i_{X_{J}^{c}}\theta)$
[…]
$=i_{X_{J}^{c}}(c\theta+i_{X_{J}^{c}}\omega) (\text{since }\mathcal{L}_{X_{J}^{c}}\theta=c\theta\mbox{ and }\omega=-d\theta)$
$=ci_{X_{J}^{c}}\theta\quad(\text{since }i_{X_{J}^{c}}i_{X_{J}^{c}}\omega=0)$
$=cJ.$
Therefore, $J$ satisfies $\mathcal{L}_{X_{J}^{c}}J=cJ$, and thus $J$ is a scaling function with parameter $c$. ∎
We now extend our analysis by exploring the relationship between a conformally Hamiltonian vector field and its associated function. In particular, we show that when the vector field satisfies a specific scaling condition on the symplectic potential, the function can be adjusted to become a scaling function:
@@
Let $(M,\omega,\theta)$ be a connected exact symplectic manifold, and let $F:M\to\mathbb{R}$ be a smooth function. Suppose $X_{F}^{c}$ is a conformally Hamiltonian vector field with parameter $c$, associated with $F$. If $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$, then there exists a function $\tilde{F}$, differing from $F$ by at most an additive constant, which is a scaling function with parameter $c$ for the vector field $X_{F}^{c}$.
@@
Suppose $\mathcal{L}_{X_{F}^{c}}\theta=c\theta$. By Lemma 5.7, we know that $F=i_{X_{F}^{c}}\theta+k_{1}$ for some constant $k_{1}$. Define $\tilde{F}=F-k_{1}$, so that $\tilde{F}=i_{X_{F}^{c}}\theta$. Clearly, $d\tilde{F}=dF=c\theta+i_{X_{F}^{c}}\omega$, and $X_{F}^{c}=X_{\tilde{F}}^{c}$.
Then, we compute:
$\mathcal{L}_{X_{\tilde{F}}^{c}}\tilde{F}$ $=i_{X_{\tilde{F}}^{c}}d\tilde{F}$
$=i_{X_{F}^{c}}(d(i_{X_{F}^{c}}\theta))\quad(\text{since }X_{F}^{c}=X_{\tilde{F}}^{c})$
[…]
@@
$=c\tilde{F}.$
Therefore, $\mathcal{L}_{X_{\tilde{F}}^{c}}\tilde{F}=c\tilde{F}$, as required. ∎
Suppose $G$ is a group acting on $Q$, with an induced cotangent lifted action on $T^{*}Q$. It is well known that the cotangent lifted action admits an equivariant momentum map, which is expressed by the formula:
$\left\langle J(z),\xi\right\rangle=\left\langle p,\xi_{Q}(q)\right\rangle,$
where $z=(q,p)\in T^{*}Q$, with $q=\pi(z)$ as the base point under the canonical projection $\pi:T^{*}Q\to Q$, and $p\in T^{*}_{q}Q$. In the Abelian case, the notion of equivariance simplifies to invariance, as the adjoint action is trivial. The equivariance of the momentum map is crucial, as it ensures that the momentum map is compatible with the group action, a necessary condition for performing symplectic reduction.
However, the conformal momentum map associated with the scaled cotangent lifted action, as described in equation (8), is not equivariant. Despite this, we have shown previously that the scaled cotangent lifted action is infinitesimally conformally invariant. As we will demonstrate below, the action is, in fact, fully conformally invariant. To establish this, we begin with a definition that extends the concept of equivariance.
@@
Let $(M,\omega,\theta)$ be an exact symplectic manifold with a conformally symplectic action $\Phi$ (with parameter $c$) of the multiplicative group $\mathbb{R}^{+}$ on $M$, and let $J$ be a conformal momentum map for $\Phi$. Then $J$ is said to be conformally equivariant if for all $g\in\mathbb{R}^{+}$,
@@
where $\Psi$ is the action of $\mathbb{R}^{+}$ on $Q$.
Since here the group is $\mathbb{R}^{+}$, the adjoint action is trivial. Hence,